OpenStudy (anonymous):

Set up and evaluate the definite integral for the area of the surface formed by revolving the graph of y=9-x^2 about the y-axis

OpenStudy (anonymous):

I know the answer is E, but how do I get that answer?

OpenStudy (tkhunny):

Please show the standard, general formulation for this sort of thing. Let's see if you're on the right page.

OpenStudy (anonymous):

Ok hold on...

OpenStudy (anonymous):

\[2\pi \int\limits_{-3}^{3}\sqrt{1+(-2y)^2} dy\] Is this the correct equation?

OpenStudy (anonymous):

Well sorta?

OpenStudy (anonymous):

The -2y; I got that from getting the derivative of 9-y^2

OpenStudy (tkhunny):

Two things funny about that. 1) Where is the missing 'x' for the radius of the pieces you are spinning? 2) Are you sure you want [-3,3] rotated through 2\(\pi\)? You may be counting things twice. What say you? It may help if you answer my question directly. The GENERAL way to proceed is: About the y-axis \(\int\limits_{a}^{b}2\pi x\sqrt{1+{\left(\dfrac{dx}{dy}\right)^{2}}}\;dy\)

OpenStudy (anonymous):

Got it! 2pi intergal min 0 max 3 and place the x in front of the square root. Everything inside the sqrt is correct

OpenStudy (anonymous):

Thank you! I tried 2pi intergal min 0 max 3 before but this whole time I was missing the x in front of the sqrt

OpenStudy (tkhunny):

If \(y = 9-x^{2}\;then\;\dfrac{dy}{dx} = -2x\), but we don't need that. We need \(\dfrac{dx}{dy} = \dfrac{1}{-2x}\)